A sphygmomanometer is used to measure the (gauge) diastolic blood pressure in a patient's veins level with their heart. This diastolic pressure is found to be 45.0 mmHg.
What is the absolute diastolic pressure in this patient's veins ? (to 3 s.f and in kPa)
(Patm = 760 mmHg = 101 kPa)
1
Expert's answer
2020-04-23T12:54:28-0400
"P=760+45=805\ mmHg"

"P=\frac{101}{760}805=107\ kPa"